OpenStudy (anonymous):

Got another i cannot seem to Algebra correctly any help is appreciated. (.3-X)^2\(.2+X)=.180

OpenStudy (stamp):

\[\frac{(0.3-x)^2}{(0.2+x)}=0.180\]\[ 0.180(0.2+x)=(0.3-x)^2\]Distribute and solve

OpenStudy (anonymous):

right, so far ive done .036+ .18x = .09 -.6x +x^2 when i solve from ehre i get \[x^2 -.6x +.09= .036 +.18x \] then -.036 -.18x from both sides give \[x^2 -.42x +.054\]

OpenStudy (anonymous):

when i apply the quadratic to this equation it gives an unreal answer for both + and - i cannot have an unreal... i have to be making a mistake but ive ran through around 4 times so far

OpenStudy (stamp):

http://www.wolframalpha.com/input/?i=%28.3-x%29^2 \%28.2%2Bx%29%3D.180

OpenStudy (phi):

-.6x - 0.18x = - 0.78x

OpenStudy (anonymous):

just got it thanks again,